OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations eliminated today? a)1300 b)1120 c)708 D)412

OpenStudy (anonymous):

x+y=1120 3x-2y=1300 take 3 times the top equation and subtract the bottom 3(x+y)- (3x-2y)=3*1120-1300 and then 3y+2y=2060 y=412 choice d

OpenStudy (anonymous):

thank u